Run Plan Updated on 13 Oct 2005 (by Dave Gill) ---------------------------------------------- NOTES: ====== 1) We MUST try to have all the TEC and M13 tune studying done before Tuesday. If all goes well this should not be a problem. 2) There are only 2 shifts available for testing if the changed production target position is an improvement. 3) The Solenoid ramp is scheduled for next Thursday to avoid being connected to the power supply when BC hydro is testing. This is what forces 2). 4) It was decided that some items would have to wait until the end of the run when we have ramped down the solenoid. Week 1 ====== 10/13/05 TEC Operating conditions: Efficiency tests Quads off tests Aperture run(s) -------------------------------------------- 10/14/05 M13 Tune: Slit Settings established in 2005 Edge scan Check 2004 tune repeatable ------------------------------------------- 10/15/05 M13 Tune: Establish tunes for polarization testing Spot size vs Q6, Q7 Position in TEC vs B2 -------------------------------------------- 10/16/05 M13 Tune: Complete tune studies Test Spare TDC (Renee): Can take from 30 minutes to several hours. Could be done while the DS scint is being installed. If the spare TDC does not work correctly, Renee proposes to remove the TEC inputs and replace them by the presently missing inputs for taking the T0 and alignment data. This would give her more time to fix the TDCs during the long maintenance on Oct 18-19th. Alignment: Install DS Scintillator trigger Usual 120 MeV/c pion data -------------------------------------------- 10/17/05 Alignment: ??? -------------------------------------------- 10/18/05 and 10/19/05 (mainetnance days) Install beam package (10/18) Total Station measurement of TEC position (10/19) -------------------------------------------------